# Date of publication: 24 April 2025

### v2

@@ 1 Introduction
Visual information has traditionally driven research in understanding human activities. Progress in computer vision techniques for tasks such as recognition and scene understanding (Kaur et al., 2024), image classification and detection (Carion et al., 2020), image segmentation (Quesada et al., 2024), video action recognition (Ma et al., 2019), and traffic surveillance (Kokilepersaud et al., 2023b) are centered around visual data. However, in real-world settings, visual processing methods face significant practical and ethical constraints, including partial occlusions (Gunasekaran and Jaiman, 2023), acquisition and environmental errors (Temel et al., 2017b, a, 2019; Prabhushankar and AlRegib, 2022), sensitivity to angles and viewpoints (Ozturk et al., 2024; Dong et al., 2022; Garg et al., 2022), privacy concerns (Mujirishvili et al., 2023; Ravi et al., 2024), and the impracticality of maintaining fixed or unobtrusive camera perspectives. In contrast, widely available wearable sensors like inertial measurement units (IMUs), electromyography (EMG), and pressure insoles capture body-centric signals unaffected by camera placement. Furthermore, they provide additional information than visual data alone.
@@
Research on human-centered applications like robotics (Chen et al., 2021), assistive technologies (Bouchabou et al., 2021), health monitoring (Wang et al., 2023), and disease diagnosis (Prabhushankar et al., 2022) utilize visual data and additional sparse data modalities. These additional modalities include skeleton data (Wang and Yan, 2023), clinical and biomarker data (Kokilepersaud et al., 2023a), and more recently language models (Damen et al., 2018) among others. However, commonly used datasets in daily activity and action understanding emphasize single-modality (Byrne et al., 2023) or scripted scenarios (Stein and McKenna, 2013) and rarely incorporate hierarchical annotations across multiple sensors (Liu et al., 2019). Addressing these gaps by designing datasets that feature continuous, unscripted recordings from both visual and non-visual sensors can enable researchers to tackle significant challenges related to privacy, multimodal learning strategies, and hierarchical activity recognition, thus promoting robust, real-world human activity understanding.
@@
Gathering continuous, real-world data across multiple environments using various sensor platforms is a significant challenge. Notably, each of the $12$ sensors used in DARai and listed in Table 3, have different sampling rates and timestamps’ protocol, making device synchronization challenging. This is addressed by implementing a multi-threaded system with a global timestamp signal to align all start and end time of data streams effectively. More information on sensor setup and calibration is provided in Appendix  LABEL:Appendix:_Sensor-Setup. We achieve a synchronization drift of 1 ms in 24-hour timespan. Depth sensors overheat during long recording sessions, reducing depth estimation quality and requiring periodic cool downs. Wearable sensors are connected wirelessly, and sometimes encounter connectivity drops. Such connectivity issues are overcome with post-recording recovery from device internal memory, thereby necessitating using sensors that provide significant internal memory. A further challenge is the large volume of data. Each participant in a session (spanning between 30 minutes to 1 hour) generates up to 300 GB of data. This necessitates local storage before transferring processed subsets to the servers. Details on data size and format are available in Appendix  LABEL:Appendix:_Structure.
@@ 3.1 Data collection process and considerations
Fifty participants (male and female) are recruited through a volunteer call and their activities are recorded in ten environments across five locations, including living spaces, home offices, and kitchens, with variations in lighting, time of day, air conditioning, and background noise. Each session includes counterfactual activity instances, where participants perform tasks with enforced variations, such as moving a heavy box instead of a light one. Some of these variations are detected by specific sensors like insole pressure sensors in Figure 1. All participants provide informed consent, are fairly compensated, and could opt out at any time, following an IRB-approved protocol that ensures participants’ privacy by removing personally identifiable information except for explicitly approved demonstration cases. Data is collected from 12 sensor devices, including Kinect, LiDAR, wearable IMUs on both hands, and stationary IMUs positioned on desks and floors in two environments, covering motion, physiological responses, and environmental conditions. Table 3 lists the sensors, their output modalities, and number of sensors used. Cameras are placed at different angles to capture standing and sitting postures, allowing natural movement rather than requiring participants to face a fixed camera, unlike other datasets (DelPreto et al., 2022; Liu et al., 2019; Qiao et al., 2022; Sener et al., 2022). After the recordings, participants complete two questionnaires: the first gathers information about their familiarity with the tasks, exhaustion levels, and BMI, while the second assesses their experience with the sensors. The survey results are provided in Appendix LABEL:Appendix:_Surveys, while additional technical details on sensor specifications, calibration, and sampling rates are in Appendix LABEL:Hardware, with a summary of sensor data modalities in Table LABEL:tab:sensors_spec.
@@ 3.2 Annotations
Some activities in the dataset do not span over all levels of the hierarchy. In Figure 3, the L2 action Rest does not contain any L3 procedures. This flexibility allows both procedural activities such as Exercise or Playing a game, and non-procedural activities including Sleeping or Watching, to fit within the same hierarchy structure and definition. In contrast, other hierarchical datasets, including (Song et al., 2024; Sener et al., 2022; Zhou et al., 2018a; Bansal et al., 2022; Kuehne et al., 2014a), have only considered procedural activities in their design. We provide the number of unique labels at each level of the hierarchy, along with examples, in Table 3. Further details about the annotation workflow are provided in Appendix LABEL:Appendix:_Annotation.
@@ 3.3 Data Preprocessing
The DARai multisensory setup records over 200 hours of data across 20 modalities from 12 sensor devices, resulting in more than 20 TB of raw recordings. Appendix LABEL:Appendix:_Structure and Table LABEL:tab:sensors_spec in Appendix LABEL:Hardware detail the specific data formats, sampling rates, and resolutions for each sensor. The raw video data is anonymized by blurring participant faces. A standard imputation method is applied for missing values on hand EMG signals, insole pressure, and biomonitoring signals along with subject-specific normalization (Gasparini et al., 2022) for each data sample. The processed data are segmented into activity-level files based on L1 annotations, making the activity samples ready for machine learning and deep learning applications. The dataset is open-sourced and available on IEEE Dataport.
@@ 4 Experiment Setup for ML Applications
In our experiments, we use a subset of the data modalities available in DARai, including RGB and Depth data from different camera views, IMU from both hands, EMG of both forearm muscles, insole pressure from feet, five biomonitoring signals, and gaze tracking data. Figure 5 summarizes the insights offered by each data modality, along with their dimensionality and resolution. The insights include motion dynamics, hand orientation, muscle activation patterns, foot pressure distribution and balance, physiological responses to varying tasks, and visual attention focus and patterns. Additional sensor data provided beyond this subset can further support research in IoT and home assistive robotics. For all experiments, we adopt a cross-subject evaluation method. Individual subjects are exclusively assigned to either the train or test set, thereby ensuring that no subject’s data is used in both training and testing. To maintain consistency across all experiments, the sets of subject IDs used for training and testing remain fixed. For reproducibility and further customization, our specific train/test subject split configuration is provided as a Python script in our GitHub repository  [^1] https://github.com/olivesgatech/DARai. For all classification tasks, we use class-wise top-1 accuracy. The class-wise accuracy metric is chosen because the hierarchical structure of human activities does not necessarily include all actions or procedures in every instance, resulting in a nonuniform distribution of lower-level classes. More details about the class distribution can be found in Appendix LABEL:Appendix:_Annotation.
We highlight the key observations from our data collection, sensor setup and each experiment in Table LABEL:tab:key_observations. A more comprehensive version of this table is provided in Appendix LABEL:Appendix:_observation-challenges.
@@ 5.1 Visual Data Benchmarking
Given the rise of large-scale pretrained models, we examined whether pretraining compensates for the diversity of daily activities. We compared models trained from scratch on DARai with models first pretrained on Kinetics 400 and then fine-tuned on DARai. For this benchmark, we evaluated R3D (Tran et al., 2018), MViT-small (Li et al., 2022), and Swin-tiny (Liu et al., 2022). Table LABEL:tab:visual-only-l1 in Appendix LABEL:Appendix:_ablation_tables compares their top-1 accuracy on combined and individual environment classes, while Table 5 reports performance by camera view at hierarchy levels L2 and L3. Although accuracy is high at the activity level (L1), all models show a notable drop at finer-grained levels (L2, L3).
@@
Analyzing these results presented in Table LABEL:tab:visual-only-l1 and Table 5 reveals a number of key observations:
@@ 6 Multimodal and Hierarchical Human Activities Benchmark
In this section, we evaluate unimodal and multimodal approaches on DARai across the three hierarchical levels (L1, L2, L3). We employ a transformer-based architecture, introduced in Appendix LABEL:Appendix:_ablation_tables, to analyze each modality alone and in pairwise combinations. Figure 8 presents these experiments results, highlighting both individual performance differences and the advantages of fusing complementary data modalities.
(a) A transformer-based architecture was trained from scratch on each modality separately for this experiment, using Swin-t for visual data modalities and the modified transformer introduced in Appendix LABEL:Appendix:_ablation_tables for other data modalities.
(b) A transformer-based architecture was trained to fuse each modality pairs separately for this experiment, using Swin-t for visual data modalities and the modified transformer introduced in Appendix LABEL:Appendix:_ablation_tables for other data modalities.
@@ 6.1 Unimodal Analysis and Comparison
Table 6 summarizes the effectiveness of different sensing modalities in recognizing human activities. Each modality excels in specific contexts but also has limitations. For example, gaze data is effective for activities involving sustained visual fixation, while EMG signals capture distinct muscle activation patterns. Biomonitoring data provide insights into physiological changes but lack fine-grained resolution for differentiating similar exertion levels. Given such complementary effectiveness, it is common to fuse these modalities to improves accuracy and robustness. A detailed study on fusion based on DARai is in Appendix LABEL:Appendix:_ablation_tables and Figure LABEL:fig:unimodal-cm.
@@ 7 Temporal Dependencies Benchmark
Understanding temporal dependencies is crucial for modeling human activities, which unfold hierarchically over time. Unlike existing datasets with predefined step boundaries, DARai presents a continuous, unstructured setting where activities transition fluidly across different granularities, posing challenges in recognizing hierarchical structure and temporal dependencies. To assess these dependencies, we focus on temporal activity localization and activity anticipation. Performance is evaluated using Mean over Classes (MoC), which measures per-class performance and averages across all classes, following the long-term action anticipation protocol (Gong et al., 2022; Abu Farha et al., 2021; Sener et al., 2020; Ke et al., 2019). For consistency, results are averaged over three different seeds. For the activity localization result, refer to LABEL:sec:app-localization.

### v3

@@ 1 Introduction
Visual information has traditionally driven research in understanding human activities. Progress in computer vision techniques for tasks such as recognition and scene understanding (Kaur et al., 2024), image classification and detection (Carion et al., 2020), image segmentation (Quesada et al., 2024), video action recognition (Ma et al., 2019), and traffic surveillance (Kokilepersaud et al., 2023c) are centered around visual data. However, in real-world settings, visual processing methods face significant practical and ethical constraints, including partial occlusions (Gunasekaran and Jaiman, 2023), acquisition and environmental errors (Temel et al., 2017b, a, 2019; Prabhushankar and AlRegib, 2022), sensitivity to angles and viewpoints (Ozturk et al., 2024; Dong et al., 2022; Garg et al., 2022), privacy concerns (Mujirishvili et al., 2023; Ravi et al., 2024), and the impracticality of maintaining fixed or unobtrusive camera perspectives. In contrast, widely available wearable sensors like inertial measurement units (IMUs), electromyography (EMG), and pressure insoles capture body-centric signals unaffected by camera placement. Furthermore, they provide additional information than visual data alone.
@@
Research on human-centered applications like robotics (Chen et al., 2021), assistive technologies (Bouchabou et al., 2021), health monitoring (Wang et al., 2023), and disease diagnosis (Prabhushankar et al., 2022) utilize visual data and additional sparse data modalities. These additional modalities include skeleton data (Wang and Yan, 2023), clinical and biomarker data (Kokilepersaud et al., 2023b), and more recently language models (Damen et al., 2018) among others. However, commonly used datasets in daily activity and action understanding emphasize single-modality (Byrne et al., 2023) or scripted scenarios (Stein and McKenna, 2013) and rarely incorporate hierarchical annotations across multiple sensors (Liu et al., 2019). Addressing these gaps by designing datasets that feature continuous, unscripted recordings from both visual and non-visual sensors can enable researchers to tackle significant challenges related to privacy, multimodal learning strategies, and hierarchical activity recognition, thus promoting robust, real-world human activity understanding.
@@
Gathering continuous, real-world data across multiple environments using various sensor platforms is a significant challenge. Notably, each of the $12$ sensors used in DARai and listed in Table 3, have different sampling rates and timestamps’ protocol, making device synchronization challenging. This is addressed by implementing a multi-threaded system with a global timestamp signal to align all start and end time of data streams effectively. More information on sensor setup and calibration is provided in Appendix  D. We achieve a synchronization drift of 1 ms in 24-hour timespan. Depth sensors overheat during long recording sessions, reducing depth estimation quality and requiring periodic cool downs. Wearable sensors are connected wirelessly, and sometimes encounter connectivity drops. Such connectivity issues are overcome with post-recording recovery from device internal memory, thereby necessitating using sensors that provide significant internal memory. A further challenge is the large volume of data. Each participant in a session (spanning between 30 minutes to 1 hour) generates up to 300 GB of data. This necessitates local storage before transferring processed subsets to the servers. Details on data size and format are available in Appendix  C.
@@ 3.1 Data collection process and considerations
Fifty participants (male and female) are recruited through a volunteer call and their activities are recorded in ten environments across five locations, including living spaces, home offices, and kitchens, with variations in lighting, time of day, air conditioning, and background noise. Each session includes counterfactual activity instances, where participants perform tasks with enforced variations, such as moving a heavy box instead of a light one. Some of these variations are detected by specific sensors like insole pressure sensors in Figure 1. All participants provide informed consent, are fairly compensated, and could opt out at any time, following an IRB-approved protocol that ensures participants’ privacy by removing personally identifiable information except for explicitly approved demonstration cases. Data is collected from 12 sensor devices, including Kinect, LiDAR, wearable IMUs on both hands, and stationary IMUs positioned on desks and floors in two environments, covering motion, physiological responses, and environmental conditions. Table 3 lists the sensors, their output modalities, and number of sensors used. Cameras are placed at different angles to capture standing and sitting postures, allowing natural movement rather than requiring participants to face a fixed camera, unlike other datasets (DelPreto et al., 2022; Liu et al., 2019; Qiao et al., 2022; Sener et al., 2022). After the recordings, participants complete two questionnaires: the first gathers information about their familiarity with the tasks, exhaustion levels, and BMI, while the second assesses their experience with the sensors. The survey results are provided in Appendix F.4, while additional technical details on sensor specifications, calibration, and sampling rates are in Appendix D.1, with a summary of sensor data modalities in Table 9.
@@ 3.2 Annotations
Some activities in the dataset do not span over all levels of the hierarchy. In Figure 3, the L2 action Rest does not contain any L3 procedures. This flexibility allows both procedural activities such as Exercise or Playing a game, and non-procedural activities including Sleeping or Watching, to fit within the same hierarchy structure and definition. In contrast, other hierarchical datasets, including (Song et al., 2024; Sener et al., 2022; Zhou et al., 2018a; Bansal et al., 2022; Kuehne et al., 2014a), have only considered procedural activities in their design. We provide the number of unique labels at each level of the hierarchy, along with examples, in Table 3. Further details about the annotation workflow are provided in Appendix E.
@@ 3.3 Data Preprocessing
The DARai multisensory setup records over 200 hours of data across 20 modalities from 12 sensor devices, resulting in more than 20 TB of raw recordings. Appendix C and Table 9 in Appendix D.1 detail the specific data formats, sampling rates, and resolutions for each sensor. The raw video data is anonymized by blurring participant faces. A standard imputation method is applied for missing values on hand EMG signals, insole pressure, and biomonitoring signals along with subject-specific normalization (Gasparini et al., 2022) for each data sample. The processed data are segmented into activity-level files based on L1 annotations, making the activity samples ready for machine learning and deep learning applications. The dataset is open-sourced and available on IEEE Dataport.
@@ 4 Experiment Setup for ML Applications
In our experiments, we use a subset of the data modalities available in DARai, including RGB and Depth data from different camera views, IMU from both hands, EMG of both forearm muscles, insole pressure from feet, five biomonitoring signals, and gaze tracking data. Figure 5 summarizes the insights offered by each data modality, along with their dimensionality and resolution. The insights include motion dynamics, hand orientation, muscle activation patterns, foot pressure distribution and balance, physiological responses to varying tasks, and visual attention focus and patterns. Additional sensor data provided beyond this subset can further support research in IoT and home assistive robotics. For all experiments, we adopt a cross-subject evaluation method. Individual subjects are exclusively assigned to either the train or test set, thereby ensuring that no subject’s data is used in both training and testing. To maintain consistency across all experiments, the sets of subject IDs used for training and testing remain fixed. For reproducibility and further customization, our specific train/test subject split configuration is provided as a Python script in our GitHub repository  [^1] https://github.com/olivesgatech/DARai. For all classification tasks, we use class-wise top-1 accuracy. The class-wise accuracy metric is chosen because the hierarchical structure of human activities does not necessarily include all actions or procedures in every instance, resulting in a nonuniform distribution of lower-level classes. More details about the class distribution can be found in Appendix E.
We highlight the key observations from our data collection, sensor setup and each experiment in Table LABEL:tab:key_observations. A more comprehensive version of this table is provided in Appendix B.
@@ 5.1 Visual Data Benchmarking
Given the rise of large-scale pretrained models, we examined whether pretraining compensates for the diversity of daily activities. We compared models trained from scratch on DARai with models first pretrained on Kinetics 400 and then fine-tuned on DARai. For this benchmark, we evaluated R3D (Tran et al., 2018), MViT-small (Li et al., 2022), and Swin-tiny (Liu et al., 2022). Table 16 in Appendix G compares their top-1 accuracy on combined and individual environment classes, while Table 5 reports performance by camera view at hierarchy levels L2 and L3. Although accuracy is high at the activity level (L1), all models show a notable drop at finer-grained levels (L2, L3).
@@
Analyzing these results presented in Table 16 and Table 5 reveals a number of key observations:
@@ 6 Multimodal and Hierarchical Human Activities Benchmark
In this section, we evaluate unimodal and multimodal approaches on DARai across the three hierarchical levels (L1, L2, L3). We employ a transformer-based architecture, introduced in Appendix G, to analyze each modality alone and in pairwise combinations. Figure 8 presents these experiments results, highlighting both individual performance differences and the advantages of fusing complementary data modalities.
(a) A transformer-based architecture was trained from scratch on each modality separately for this experiment, using Swin-t for visual data modalities and the modified transformer introduced in Appendix G for other data modalities.
(b) A transformer-based architecture was trained to fuse each modality pairs separately for this experiment, using Swin-t for visual data modalities and the modified transformer introduced in Appendix G for other data modalities.
@@ 6.1 Unimodal Analysis and Comparison
Table 6 summarizes the effectiveness of different sensing modalities in recognizing human activities. Each modality excels in specific contexts but also has limitations. For example, gaze data is effective for activities involving sustained visual fixation, while EMG signals capture distinct muscle activation patterns. Biomonitoring data provide insights into physiological changes but lack fine-grained resolution for differentiating similar exertion levels. Given such complementary effectiveness, it is common to fuse these modalities to improves accuracy and robustness. A detailed study on fusion based on DARai is in Appendix G and Figure 29.
@@ 7 Temporal Dependencies Benchmark
Understanding temporal dependencies is crucial for modeling human activities, which unfold hierarchically over time. Unlike existing datasets with predefined step boundaries, DARai presents a continuous, unstructured setting where activities transition fluidly across different granularities, posing challenges in recognizing hierarchical structure and temporal dependencies. To assess these dependencies, we focus on temporal activity localization and activity anticipation. Performance is evaluated using Mean over Classes (MoC), which measures per-class performance and averages across all classes, following the long-term action anticipation protocol (Gong et al., 2022; Abu Farha et al., 2021; Sener et al., 2020; Ke et al., 2019). For consistency, results are averaged over three different seeds. For the activity localization result, refer to G.2.